OpenStudy (anonymous):

HELP: Solve: 35q^2+33q+4=0

OpenStudy (anonymous):

@dpasingh

OpenStudy (cwrw238):

you need to check out factors of 35 and + 4 and figure out how you can get middle coefficient +33 35 = 5 * 7 4 = 4 * 1 28 + 5 = 33 thats it

OpenStudy (cwrw238):

7*4 + 5*1

OpenStudy (cwrw238):

giving (5q+ 4 )(7q + 1 )

OpenStudy (cwrw238):

(5q + 4)(7q + 1) = 0

OpenStudy (anonymous):

\[\huge 35q^2+33q+4=0\] \[\huge 35q^2+28q + 5q+4=0\] \[\huge 7q(5q+4) +1( 5q+4)=0\] \[\huge (7q +1)( 5q+4)=0\] (7q +1)=0 or ( 5q+4)=0 q= -1/7 , -4/5

OpenStudy (anonymous):

thank you both.

OpenStudy (cwrw238):

yw the middle coefficient is obtained as in diagram